# On pseudo-effective cones of projective bundles and volume function

### v3

@@ 6.1. Harder-Narasimhan filtration of tensor product
Then $\Im_{A}=\Im_{\overline{A}}$. Also for any two subsets $A_{1}\subseteq A_{2}\subseteq\Theta$, we have $\Im_{A_{1}}\subseteq\Im_{A_{2}}$. For any $\alpha=(a,b)\in\Theta$, we define
@@
Locally in Zariski topology, the Harder-Narsimhan filtrations split. Note that using this splitting, one can show that locally we have the required isomorphism of the form (9). These natural isomorphisms do not depend on the choice of the splitting, and therefore glue together to give the global isomorphism $\Im_{A}/\Im_{A^{\prime}}\cong\bigoplus\limits_{\alpha\in A^{\prime\prime}}\mathcal{Q}_{\alpha}$. ∎
@@ 6.2. Harder Narasimhan Filtration for $\operatorname{{Sym}}^{mn}(E)$ :
For $\alpha=(a_{i})_{i=1}^{mn}\in\Phi$, we define $E_{\alpha}=\bigotimes\limits_{i=1}^{mn}E_{a_{i}}$. For a subset $A\subseteq\Phi$, we define $E_{A}:=\bigoplus\limits_{\alpha\in A}E_{\alpha}.$
@@ 6.3. Harder Narasimhan Filtration for $\operatorname{{Sym}}^{ln}(F)$ :
Let $\Phi^{\prime}=\{0,1,2,\cdots,h-1\}^{ln}$. For $\beta=(b_{i})_{i=1}^{ln}\in\Phi^{\prime}$, we define $F_{\beta}:=\bigotimes\limits_{i=1}^{ln}F_{b_{i}}$, and for a subset $B\subseteq\Phi^{\prime}$, we define $F_{B}:=\bigoplus\limits_{\beta\in B}F_{\beta}$.
@@ 7. Acknowledgement
We are grateful to the referee for a careful reading and the many suggestions which improved the paper. The first author is supported financially by SERB-NPDF fellowship (File no : PDF/2021/00028). The first author also like to thank IIT Bombay for its hospitality where the work is initiated.

### v4

@@ 6.1. Harder-Narasimhan filtration of tensor product
i.e. $\Im_{A}$ is the sum of subbundles generated by $\Bigl{\{}\Im_{\alpha}:\alpha\in A\Bigr{\}}.$ Write by convention $E_{\emptyset}=0$. Then $\Im_{A}=\Im_{\overline{A}}$. Also for any two subsets $A_{1}\subseteq A_{2}\subseteq\Theta$, we have
$\Im_{A_{1}}\subseteq\Im_{A_{2}}.$
For any $\alpha=(a,b)\in\Theta$, we define
@@
Let $\alpha\in A^{\prime}$ , $\beta\in\Theta$, and $\beta\geq\alpha$. As $A$ is saturated, we have $\beta\in A$. But $\beta$ is not a minimal element of $A$. Thus $\beta\in A^{\prime}$. This shows that $A^{\prime}$ is saturated.
Let $A=\bigl{\{}\alpha_{1},\alpha_{2},\cdots,\alpha_{n}\bigr{\}}$ such that $A^{\prime\prime}=\bigl{\{}\alpha_{1},\alpha_{2},\cdots,\alpha_{m}\bigr{\}}$ and $A^{\prime}=\bigl{\{}\alpha_{m+1},\alpha_{m+2},\cdots,\alpha_{n}\bigr{\}}$.
Recall that for any $\alpha=(a,b)\in A$, we define $\Im_{\alpha}=E_{a}\otimes F_{b}$ and $\mathcal{Q}_{\alpha}=Q_{a}\otimes Q^{\prime}_{b}$. We note that there is a surjection
$\psi_{\alpha}:\Im_{\alpha}=E_{a}\otimes F_{b}\longrightarrow Q_{a}\otimes F_{b}\longrightarrow\mathcal{Q}_{\alpha}=Q_{a}\otimes Q^{\prime}_{b}\longrightarrow 0.$
Define
$\Psi:\Im_{A}=\sum\limits_{\alpha\in A}\Im_{\alpha}\longrightarrow\bigoplus\limits_{\alpha\in A^{\prime\prime}}\mathcal{Q}_{\alpha}$
as follows :
$\Psi(\sum\limits_{j=1}^{n}x_{j})=\bigl{(}\psi_{\alpha_{1}}(x_{1}),\psi_{\alpha_{2}}(x_{2}),\cdots,\psi_{\alpha_{m}}(x_{m})\bigr{)},$
where $x_{j}\in\Im_{\alpha_{j}}$ for each $j\in\{1,2,\cdots,n\}$. We fix the notation $\alpha_{j}=(a_{j},b_{j})$ for $j$ satisfying $1\leq j\leq n$. Note that if $\alpha\leq\beta$ for some $\alpha,\beta\in A$, then we have $\Im_{\beta}\hookrightarrow\Im_{\alpha}$. Moreover, if $\alpha\leq\beta$ for some $\alpha,\beta\in A$ and $\alpha\neq\beta$, then $\Im_{\beta}\subseteq\ker(\psi_{\alpha})$.
First we claim the following : if $\sum\limits_{j=1}^{n}x_{j}=0\in\Im_{A}$, then
$\Psi(\sum\limits_{j=1}^{n}x_{j})=\bigl{(}\psi_{\alpha_{1}}(x_{1}),\psi_{\alpha_{2}}(x_{2}),\cdots,\psi_{\alpha_{m}}(x_{m})\bigr{)}=(\underbrace{0,0,0,\cdots,0}_{m-times})\in\bigoplus\limits_{\alpha\in A^{\prime\prime}}\mathcal{Q}_{\alpha}.$
Now suppose $\psi_{\alpha_{i}}(x_{i})\neq 0$ for some $\alpha_{i}$ with $1\leq i\leq m$. Consider the set
$T:=\Bigl{\{}l\mid 1\leq l\leq n,(a_{l},b_{l})\geq(a_{i},b_{i})$ and $(a_{l},b_{l})\neq(a_{i},b_{i})\Bigr{\}}.$
We then have
$0=\sum\limits_{j=1}^{n}x_{j}=x_{i}+\sum\limits_{l\in T}x_{l}+\sum\limits_{l\notin T,l\neq i}x_{l}$
Note that for any $l\in T$, we have $\Im_{(a_{l},b_{l})}\subseteq\ker(\psi_{\alpha_{i}})$. Thus $\sum\limits_{l\in T}x_{l}\in\sum\limits_{l\in T}\Im_{(a_{l},b_{l})}\subseteq\ker(\psi_{\alpha_{i}}).$
Therefore
$x_{i}+\sum\limits_{l\notin T,l\neq i}x_{l}\in\ker(\psi_{\alpha_{i}})\subseteq\Im_{\alpha_{i}}.$
We will now show that $\sum\limits_{l\notin T,l\neq i}x_{l}\in\ker(\psi_{\alpha_{i}})$ and this will contradict the fact that $x_{i}\notin\ker(\psi_{\alpha_{i}}).$
Note that
$\sum\limits_{l\notin T,l\neq i}x_{l}\in\Im_{(a_{i},b_{i})}.$
Also
$\sum\limits_{l\notin T,l\neq i}x_{l}\in\sum\limits_{l\in T,l\neq i}\Im_{(a_{l},b_{l})}\subseteq\Im_{(s,t)},$
where $s=\min\{a_{l}\mid 1\leq l\leq n,l\notin T,l\neq i\}$ and $t=\min\{b_{l}\mid 1\leq l\leq n,l\notin T,l\neq i\}.$ Let $s=a_{p}$ for some $p\notin T$ and $p\neq i$, and $t=b_{q}$ for some $q\notin T$ and $q\neq i$.
Then we observe that
$\sum\limits_{l\notin T,l\neq i}x_{l}\in\Im_{(a_{i},b_{i})}\cap\Im_{(a_{p},b_{q})}=\Im_{(\max\{a_{p},a_{i}\},\max\{b_{q},b_{i}\})}.$
Next we show that $(\max\{a_{p},a_{i}\},\max\{b_{q},b_{i}\})\neq(a_{i},b_{i})$. If not, then $\max\{a_{p},a_{i}\}=a_{i}$ and $\max\{b_{q},b_{i}\}=b_{i}$. This implies $a_{i}\geq a_{p}$ and $b_{i}\geq b_{q}$. Since $p\notin T$ and $p\neq i$, we have either $a_{p}<a_{i}$ or $b_{p}<b_{i}$. Consider the following two cases :
  1. (1)
    Case 1 : If $a_{p}<a_{i}$, then we arrive at a contradiction as $a_{i}\geq a_{p}.$
  2. (2)
    Case 2 : If $b_{p}<b_{i}$, then by definition of $b_{q}$ we have $b_{q}\leq b_{p}<b_{i}$, and this also gives a contradiction.
Therefore $(\max\{a_{p},a_{i}\},\max\{b_{q},b_{i}\})\geq(a_{i},b_{i})$ and $(\max\{a_{p},a_{i}\},\max\{b_{q},b_{i}\})\neq(a_{i},b_{i}).$
Thus we conclude
$\sum\limits_{l\notin T,l\neq i}x_{l}\in\Im_{(\max\{a_{p},a_{i}\},\max\{b_{q},b_{i}\})}\subseteq\ker(\psi_{\alpha_{i}}).$
This completes the proof of our claim.
Now we show that $\Psi$ is well-defined. Let $\sum\limits_{j=1}^{n}x_{j}$ and $\sum\limits_{j=1}^{n}x^{\prime}_{j}\in\Im_{A}$ such that $\sum\limits_{j=1}^{n}x_{j}=\sum\limits_{j=1}^{n}x^{\prime}_{j}$. Then
$\sum\limits_{j=1}^{n}\bigl{(}x_{j}-x^{\prime}_{j})=\underbrace{0+0+0+\cdots+0}_{n-times}\in\Im_{A}.$
Hence by the previous observation, we get $\psi_{\alpha_{i}}(x_{j}-x^{\prime}_{j})=0$ for each $i$ and $j$ with $1\leq i,j\leq m$. As each $\psi_{\alpha}$ is a homomorphism, we have $\psi_{\alpha_{i}}(x_{j})=\psi_{\alpha_{i}}(x^{\prime}_{j})$ for each $i$ and $j$ satisfying $1\leq i,j\leq m$. Therefore we conclude $\Psi\bigl{(}\sum\limits_{j=1}^{n}x_{j}\bigr{)}=\Psi\bigl{(}\sum\limits_{j=1}^{n}x^{\prime}_{j}\bigr{)}.$ This proves the well-definedness of the map $\Psi$.
Clearly, we have $\Im_{A^{\prime}}\subseteq$ ker$(\Psi)$. Also, the map $\Psi$ is surjective as each $\psi_{\alpha}$ is surjective for any $\alpha\in A^{{}^{\prime\prime}}$. Next we will show that ker$(\Psi)\subseteq\Im_{A^{\prime}}$.
Let $\alpha=(a,b)\in A^{\prime\prime}$ be a minimal element and $x\in$ Ker$(\psi_{\alpha})$. For any minimal element $\alpha=(a,b)\in A^{{}^{\prime\prime}}$, we have the following commutative diagram of short exact sequences:
[…]
Note that the map $\psi_{\alpha}=\gamma_{2}\circ\theta_{4}$. Let $x\in$ Ker$(\psi_{\alpha})$. Then $(\gamma_{2}\circ\theta_{4})(x)=\gamma_{2}(\theta_{4}(x))=0$. So there exists $y\in Q_{a}\otimes F_{b+1}$ such that $\gamma_{1}(y)=\theta_{4}(x)$. Now there exists $z\in E_{a}\otimes F_{b+1}$ such that $\theta_{2}(z)=y$. This implies $\gamma_{1}(\theta_{2}(z))=\theta_{4}(x)$. By the commutativity of the diagram, we have $\gamma_{1}(\theta_{2}(z))=\theta_{4}(\beta_{1}(z))=\theta_{4}(x)$. This implies $x-\beta_{1}(z)\in$ Ker$(\theta_{4})$. Thus there exists $t\in E_{a+1}\otimes F_{b}$ such that $x-\beta_{1}(z)=\theta_{3}(t)$. In other words, $x=\beta_{1}(z)+\theta_{3}(t)$. This shows that Ker$(\psi_{\alpha})\subseteq E_{a+1}\otimes F_{b}+E_{a}\otimes F_{b+1}\subseteq\Im_{A^{\prime}}$.
Let $\sum\limits_{i=1}^{n}x_{i}\in$ Ker$(\Psi)$. Thus $\psi_{\alpha_{i}}(x_{i})=0$ for every $i\in\{1,2,\cdots,m\}$, i.e. $x_{i}\in$ Ker$(\psi_{\alpha_{i}}).$ Hence by the previous observation we have $\sum\limits_{i=1}^{n}x_{i}\in\Im_{A^{\prime}}$. Therefore we conclude that Ker$(\Psi)\subseteq\Im_{A^{\prime}}$.
We then have the required isomorphism :
$\Psi:\Im_{A}/\Im_{A^{\prime}}\cong\bigoplus\limits_{\alpha\in A^{\prime\prime}}\mathcal{Q}_{\alpha}$
@@ 6.2. Harder Narasimhan Filtration for $\operatorname{{Sym}}^{mn}(E)$ :
For $\alpha=(a_{i})_{i=1}^{mn}\in\Phi$, we define $E_{\alpha}=\bigotimes\limits_{i=1}^{mn}E_{a_{i}}$. For a subset $A\subseteq\Phi$, we define $E_{A}:=\sum\limits_{\alpha\in A}E_{\alpha}.$
@@ 6.3. Harder Narasimhan Filtration for $\operatorname{{Sym}}^{ln}(F)$ :
Let $\Phi^{\prime}=\{0,1,2,\cdots,h-1\}^{ln}$. For $\beta=(b_{i})_{i=1}^{ln}\in\Phi^{\prime}$, we define $F_{\beta}:=\bigotimes\limits_{i=1}^{ln}F_{b_{i}}$, and for a subset $B\subseteq\Phi^{\prime}$, we define $F_{B}:=\sum\limits_{\beta\in B}F_{\beta}$.
@@ 7. Acknowledgement
We are grateful to the referee for a careful reading and the many suggestions which improved the paper. The first author is partially supported by SERB-NPDF fellowship (File no : PDF/2021/00028). The second author is partially supported by SERB SRG Grant SRG/2023/001006. The first author also like to thank IIT Bombay for its hospitality where the work is initiated.